# Thermodynamic geometry in hadron resonance gas model at real and imaginary baryon chemical potential and a simple sufficient condition for quark deconfinement

### v2

@@ Abstract
The thermodynamic geometry of the hadron resonance gas model with (without) excluded volume effects (EVE) of baryons is investigated. The case with imaginary $\mu$, where $\mu$ is the baryon chemical potential, is investigated as well as the one with real $\mu$. We calculate the scalar curvature $R$ and use the $R=0$ criterion to investigate the phase structure in the $\mu^{2}$-$T$ plane where $T$ is the temperature. The curve on which $R=0$ continues analytically from the imaginary $\mu$ region, where the lattice QCD is feasible, to the real $\mu$ one. In the presence of EVE, there are rich phase structures in the large real $\mu$ region as well as the Roberge-Weiss like region where $\mu$ is imaginary and a singularity appears, while there is no phase structure in the large real $\mu$ region in the absence of EVE. The limitation temperature of the baryon gas is also obtained by using the baryon number fluctuation. The LQCD predicted critical point locates almost on the curve of the limitation temperature we determined. A simple sufficient condition, $n_{\rm B}>1/(2v_{\rm B})$, is obtained for the quark deconfinement in the large real $\mu$ region, where $n_{\rm B}$ and $v_{\rm B}$ are the net baryon number density and the volume of a baryon, respectively.
@@ I Introduction
Determination of the phase diagram of quantum chromodynamics (QCD) is an important subject not only in nuclear and particle physics but also in cosmology and astrophysics; see, e.g., Ref. Fukushima and Hatsuda (2011) and references therein. However, when the baryon number chemical potential is finite and real, the first principle calculation, namely, the lattice QCD (LQCD) simulation, is not feasible due to the infamous sign problem; see Refs. de Forcrand (2009); Nagata (2020); *Nagata:2021ugx as an example. To circumvent the sign problem, several methods are proposed and investigated, although, at present, these methods are not complete and we do not have adequate information on the equation of state (EoS) at finite baryon density.
So far, the properties of the nuclear/quark matter at high baryon density are mainly investigated by using the effective model of QCD. It is known that the LQCD results at $\mu=0$ are in good agreement with those obtained by the hadron resonance gas (HRG) model when temperature $T$ is not so large. Usually, the ideal gas approximation is used for the calculations in the HRG model. However, it is expected that repulsive effects among baryons are important at high density. If the repulsion is absent, baryon matter is realized at a sufficiently large baryon density Cleymans et al. (1986a). One of the traditional treatments for such repulsion is to consider excluded volume effects (EVE) among baryons Cleymans et al. (1986b); Kouno and Takagi (1989); Rischke et al. (1991). EVE successfully prevents baryon matter from realizing at large baryon chemical potential Cleymans et al. (1986b); for the recent review, see, e.g., Ref. Fujimoto et al. (2022) and references therein. The availability of the HRG model with EVE may be checked by using the LQCD results at finite imaginary $\mu$.
When the baryon (quark) chemical potential $\mu$ ($\mu_{\rm q}$) is pure imaginary, there is no sign problem and one can perform LQCD simulations; see Refs. de Forcrand and Philipsen (2002a); D’Elia and Lombardo (2003a, 2004); Chen and Luo (2005); D’Elia and Sanfilippo (2009) as an example. The grand canonical QCD partition function $Z(\theta)$ with pure imaginary quark chemical potential ($\mu_{\rm q}={\mu/3}=i\theta_{\rm q}T$) has the Roberge-Weiss (RW) periodicity Roberge and Weiss (1986) as
@@
Figure 1 shows the $\theta$-dependence of the net baryon density $n_{\rm B}$ at $T=T_{\rm RWL}$. In the case of the HRG model with EVE, $n_{\rm B}$ has a discontinuous singularity at $\theta=\pm\pi$, while it is continuous function of $\theta$ there in the absence of the EVE. Figure 2 shows the real $\mu$-dependence of the net baryon density $n_{\rm B}$ at $T=T_{\rm RWL}$. In the case of the HRG model with EVE, $n_{\rm B}$ saturates to the constant value $\pm 1/v_{\rm B}$ in the limit $\mu\to\pm\infty$ where $v_{\rm B}$ is a volume of a baryon, while the absolute value of it increases rapidly in the absence of the EVE as $|\mu|$ increases. It is clear that the saturation in the real $\mu$ region is strongly correlated with the singularity in the imaginary $\mu$ region. The singular behavior in the vicinity of the RWL point has much information about the nuclear EOS at high baryon density. However, the nature of the interactions at the RWL point is not clear yet.
@@
One of the powerful tools to investigate the features of the interaction in the thermodynamic system is the thermodynamic geometry; for a review, see, e.g, Ref. Ruppeiner (1995) and references therein. In the thermodynamic geometry, the scalar curvature $R$ is an important quantity. In particle and nuclear physics, Castorina, Imbrosciano and Lanteri used the thermodynamic geometry to investigate the QCD phase diagram in $\mu$-$T$ plane Castorina et al. (2018). Since there is a sign problem in finite density LQCD, they used the power series expansion method to estimate the scalar curvature $R$. They used the $R=0$ criterion to determine the pseudo-critical temperature of the crossover transition. The obtained pseudo-critical temperature is consistent with that obtained by using the standard criterion in the LQCD. They also used the HRG model without EVE to calculate $R$. They found that the pseudo-critical temperature obtained in the HRG model is somewhat higher than that in LQCD. In Ref. Castorina et al. (2019), they did the similar analyses but used also the HRG model with EVE. In this case, they found that the pseudo-critical temperature obtained by the HRG model is consistent with that in LQCD. The excluded volume effects is important for reproducing the pseudo-critical temperature.
In this paper, to investigate the nature of the EVE effect in detail, we study the HRG model with/without EVE at imaginary $\mu$ as well as real $\mu$ using the thermodynamic geometry. This paper is organized as follows. In Sec. II, the RW periodicity and transition are briefly reviewed. In Sec. III, we show our formulation of the HRG model with EVE. In Sec. IV, we briefly review the thermodynamic geometry. In Sec. V, the numerical results of the scalar curvature and the phase structures are shown. In Sec. VI, we investigate the limitation temperature of our HRG model with EVE. Using the limitation temperature, an approximate location of the critical point and a simple sufficient condition for quark deconfinement are obtained. Section VII is devoted to the summary and discussions.
@@ II Roberge-Weiss periodicity and transition
${\cal L}(\theta_{\rm q})$ $=\bar{\psi}(\gamma_{\mu}D_{\mu}-m_{0})\psi-{1\over{4}}F_{\mu\nu}^{2}-i{\theta_{\rm q}\over{\beta}}\bar{\psi}\gamma_{4}\psi,$ (5)
is the action. Here $\psi$, $A_{\mu}$, $F_{\mu\nu}$, $D_{\mu}$, and $m_{0}$ are the quark field, the gluon field, the gluon field strength, the covariant derivative and the current quark mass matrix, respectively, and $\beta=1/T$. We use the Euclidean notation in Eqs. (3) $\sim$ (5).
@@
Next, we perform another transformation of the quark and gluon fields as follows.
@@
When $T<T_{\rm RW}$ where $T_{\rm RW}$ is the RW temperature, the net quark number density $n_{\rm Q}$ is a smooth function of $\theta_{\rm q}$. However, when $T>T_{\rm RW}$, it is discontinuous at $\theta_{\rm q}=(2k+1)\pi/3$ where $k$ is any integer, due to the degeneracy of the ground state. This discontinuity is called the RW transition. Similarly, the pressure $P$ of the system is a smooth function of $\theta_{\rm q}$ when $T<T_{\rm RW}$, but it has a cusp at $\theta_{\rm q}=(2k+1)\pi/3$ when $T>T_{\rm RW}$. The RW periodicity and the RW transition are confirmed by LQCD simulations de Forcrand and Philipsen (2002b); D’Elia and Lombardo (2003b) and $T_{\rm RW}$ is estimated as $195\sim 208$ MeV for the 2+1 flavor LQCD simulation Bonati et al. (2016); Cuteri et al. (2022); Bonati et al. (2019).
Fig. 3 shows the $\theta_{\rm q}$-dependence of the net baryon number density $n_{\rm Q}/3$ obtained by the PNJL model Fukushima (2004); Ratti et al. (2006); Ghosh et al. (2006); Megias et al. (2006); Rößner et al. (2007); Kashiwa et al. (2008), when $T=T_{\rm RWL}=0.2103$ GeV. The PNJL model is one of the most successful effective models of QCD and can also reproduce several important features of QCD in the region of the imaginary chemical potential when $T$ is not so small; for example, see Ref. Kashiwa (2019) as a review. In Fig. 3, we see that the net baryon number density is discontinuous at $\theta=\pm\pi/3$. (Note that, in this calculation, $T=T_{\rm RWL}=0.2103$ GeV is larger than $T_{\rm RW}=0.201~{\rm GeV}$. Also note that the net baryon (quark) number density is pure imaginary in the region of the imaginary $\mu~(\mu_{q})$. )
@@ III Hadron resonance gas model with excluded volume effects
where $P_{\rm M}$ is the pressure of mesons. For the mesons, we use a ideal Bose gas approximation. The other thermodynamic quantities are also calculated by using the thermodynamic relations. Our formalism of the HRG model with EVE is different from that of Refs. Taradiy et al. (2019); Savchuk et al. (2020); Vovchenko et al. (2017) . Comparison between two models is shown in Refs. Oshima et al. (2025) and our model are consistent with the model in Ref. Taradiy et al. (2019); Savchuk et al. (2020); Vovchenko et al. (2017) at least below $T_{\rm RWL}$.
@@
Hence, $n_{\rm B}$ is divergent at $\theta=(2n+1)\pi$ for any integer $n$ as is seen in Fig. 1. This singularity is nothing but the RW-like singularity which occurs at $T=T_{\rm RWL}=0.2103$ GeV. The value 0.2103 GeV is slightly smaller than the value 0.2108 GeV in Ref. Oshima et al. (2025), since we have updated the data of hadron resonances Navas and others (2024). On the other hand, as was already shown in Ref. Oshima et al. (2025), the use of the Boltzmann distribution approximation hardly changes the value of $T_{\rm RWL}$. If we use the nucleon gas model instead of the baryon resonances gas model, we obtain $T_{\rm RWL}=0.316$ GeV which is much higher than $T_{\rm RW}\sim 0.2$ GeV. The effects of the baryon resonances are much important for the realization $T_{\rm RWL}\sim T_{\rm RW}$.
@@
Comparing Fig. 1 (Fig. 6) with Fig. 2 (Fig. 5), we see that the saturation (linearization) of $n_{\mathrm{B}}$ ($P$) in the limit $\mu\to\pm\infty$ is related to the RW-like singularity at $\theta=(2n+1)\pi$. This indicates that the properties of the RW-like singularity in the imaginary chemical potential region have important information of the interaction among baryons in the large real chemical potential region.
@@
It seems that $T_{\rm RWL}$ is slightly higher than $T_{\rm RW}$. Hence, our scenario of the hadron-quark transition at $\theta=(2n+1)\pi$ is as follows. As $T$ approaches $T_{\rm RWL}$, the baryons become unstable and transit to quark matter when $T$ is equal to the temperature which is slightly lower than $T_{\rm RWL}$. This transition temperature is nothing but the Roberge-Weiss temperature $T_{\rm RW}$. However, the nature of the interaction which induces the instability is not clear yet. Hence, in this paper, we will investigate the nature of the interaction using the thermodynamic geometry.
@@ IV Thermodynamic geometry
The scalar curvature $R$ can be used to investigate the nature of the interaction. It is expected that the interaction in the system is repulsive (attractive), when $R>0$ ($R<0$). Using this criterion, it can be concluded that the ideal fermion gas has the repulsive nature, while the ideal boson gas does the attractive one Janyszek and Mrugaa (1990). Hence, the interaction in baryon (meson) gas is expected to be repulsive (attractive).
@@
The thermodynamic geometry can be also used to determine the phase diagram. In $\mu$-$T$ plane, the curve on which $R=0$ may be the transition line of the phase transition. Castorina, Imbrosciano and Lanteri used the thermodynamic geometry to investigate the QCD phase diagram in $\mu$-$T$ plane Castorina et al. (2018). Since there is a sign problem in finite density LQCD, they used the power series expansion method to estimate $\phi$ and calculated the scalar curvature $R$. They used the $R=0$ criterion to determine the pseudo-critical temperature of the crossover transition. The obtained pseudo-critical temperature is consistent with that obtained by using the standard criterion in the LQCD. They also used the HRG model to estimate $\phi$ and calculated $R$. They found that the pseudo-critical temperature obtained in the HRG model is somewhat higher than that in LQCD. In Ref. Castorina et al. (2019), they did the similar analyses but used also the HRG model with EVE. In this case, they found that the pseudo-critical temperature obtained by the HRG model with EVE is consistent with that in LQCD.
@@
The scalar curvature $R$ has the same dimension as the spatial volume. Hence, it is expected that, in proximity of a second order phase transition, $R$ is proportional to $\xi^{3}$ where $\xi$ is the correlation length and diverges when the second order phase transition takes happen. Castorina, Lanteri and Mancani  Castorina et al. (2020a) studied the thermodynamic geometry in Nambu–Jona-Lasinio model. They found that the scalar curvature $R$ shows the divergent behavior at the critical temperature of the chiral phase transition when the current quark mass is zero. They also found that $R$ becomes large when the system approaches critical point when the current quark mass is finite. Zhang, Wan and Ruggieri did the similar studies by using the Quark-meson model Zhang et al. (2020). Castorina, Lanteri and Ruggieri studied the fluctuations and the thermodynamic geometry of the chiral phase transition in the Quark-meson model Castorina et al. (2020b). Murgana, Greco, Ruggieri and Zappalà studied the thermodynamic geometry of the Quark-meson model using the functional renormalization group Murgana et al. (2024).
@@
At a glance, the two criterions, namely, the $R=0$ criterion and the $|R|=\infty$ criterion, are contradicted each other. However, in many cases, the divergence of $R$ is accompanied by the change of sign of $R$ Murgana et al. (2024). In particular, the sign changing of $R$ takes place in the vicinity of the critical point where $R$ shows the divergent behavior Murgana et al. (2024). As will be seen later, there is no divergence of $R$ in the HRG model when the chemical potential $\mu$ is real. Hence, we mainly use the $R=0$ criterion to investigate the phase structure in the HRG model as was in the previous studies Castorina et al. (2018, 2019).

### v3

@@ Abstract
The thermodynamic geometry of the hadron resonance gas model with (without) excluded volume effects (EVE) of baryons is investigated. The case with imaginary $\mu$, where $\mu$ is the baryon chemical potential, is investigated as well as the one with real $\mu$. We calculate the scalar curvature $R$ and use the $R=0$ criterion to investigate the phase structure in the $\mu^{2}$-$T$ plane where $T$ is the temperature. The curve on which $R=0$ continues analytically from the imaginary $\mu$ region, where the lattice QCD is feasible, to the real $\mu$ one. In the presence of EVE, there are rich phase structures in the large real $\mu$ region as well as the Roberge-Weiss like region where $\mu$ is imaginary and a singularity appears, while there is no phase structure in the large real $\mu$ region in the absence of EVE. The limiting temperature of the baryon gas is also obtained by using the baryon number fluctuation. The LQCD predicted critical point locates almost on the curve of the limiting temperature we determined. A simple empiric sufficient condition, $n_{\rm B}>1/(2v_{\rm B})$, is obtained for the quark deconfinement in the large real $\mu$ region, where $n_{\rm B}$ and $v_{\rm B}$ are the net baryon number density and the volume of a baryon, respectively.
@@ I Introduction
The determination of the phase diagram of quantum chromodynamics (QCD) is an important subject not only in nuclear and particle physics but also in cosmology and astrophysics; see, e.g., Ref. Fukushima and Hatsuda (2011) and references therein. However, when the baryon number chemical potential is finite and real, the first-principle calculation, namely, the lattice QCD (LQCD) simulation, is not feasible due to the infamous sign problem; see Refs. de Forcrand (2009); Nagata (2020); *Nagata:2021ugx as an example. To circumvent the sign problem, several methods have been proposed and investigated, although, at present, these methods are not complete and we do not have adequate information on the equation of state (EOS) at finite baryon density.
So far, the properties of the nuclear/quark matter at high baryon density are mainly investigated by using effective models of QCD. It is known that the LQCD results at $\mu=0$ are in good agreement with those obtained by the hadron resonance gas (HRG) model when temperature $T$ is not so large. Usually, the ideal gas approximation is used for the calculations in the HRG model. However, it is expected that repulsive effects among baryons are important at high density. If the repulsion is absent, baryon matter is realized at a sufficiently large baryon density Cleymans et al. (1986a). One of the traditional treatments for such repulsion is to consider excluded volume effects (EVE) among baryons Cleymans et al. (1986b); Kouno and Takagi (1989); Rischke et al. (1991). EVE successfully prevents the realization of baryon matter at large baryon chemical potential Cleymans et al. (1986b); for the recent review, see, e.g., Ref. Fujimoto et al. (2022) and references therein. The availability of the HRG model with EVE may be checked by using the LQCD results at finite imaginary $\mu$.
When the baryon (quark) chemical potential $\mu$ ($\mu_{\rm q}$) is pure imaginary, no sign problem occurs and one can perform LQCD simulations; see Refs. de Forcrand and Philipsen (2002a); D’Elia and Lombardo (2003a, 2004); Chen and Luo (2005); D’Elia and Sanfilippo (2009) as an example. The grand canonical QCD partition function $Z(\theta)$ with pure imaginary quark chemical potential ($\mu_{\rm q}={\mu/3}=i\theta_{\rm q}T$) has the Roberge-Weiss (RW) periodicity Roberge and Weiss (1986) as
@@
Figure 1 shows the $\theta$-dependence of the net baryon density $n_{\rm B}$ at $T=T_{\rm RWL}$. In the case of the HRG model with EVE, $n_{\rm B}$ presents a discontinuous singularity at $\theta=\pm\pi$, while it is smooth function of $\theta$ at those points in the absence of the EVE. Figure 2 shows the -$\mu$ dependence of the net baryon density $n_{\rm B}$ at $T=T_{\rm RWL}$. In the case of the HRG model with EVE, $n_{\rm B}$ saturates to the constant value $\pm 1/v_{\rm B}$ in the limit $\mu\to\pm\infty$ where $v_{\rm B}$ is the volume of a baryon, while the absolute value of it increases rapidly in the absence of the EVE as $|\mu|$ increases. It is clear that the saturation in the real $\mu$ region is strongly correlated with the singularity in the imaginary $\mu$ region. The singular behavior in the vicinity of the RWL point contains significant information about the nuclear equation of state at high baryon density. However, the nature of the interactions at the RWL point is not clear yet.
@@
One of the powerful tools to investigate the features of the interaction in the thermodynamic system is the thermodynamic geometry; for a review, see, e.g, Ref. Ruppeiner (1995) and references therein. In the thermodynamic geometry, the scalar curvature $R$ is an important quantity. In particle and nuclear physics, thermodynamic geometry was used to investigate the QCD phase diagram in $\mu$-$T$ plane Castorina et al. (2018). It was found that the phase diagram obtained by the thermodynamic geometry in the HRG model with EVE is consist with the pseudo-critical temperature obtained in LQCD Castorina et al. (2019). The excluded volume effects are important for reproducing the pseudo-critical temperature.
In this paper, to investigate the nature of the EVE in detail, we study the HRG model with/without EVE at imaginary $\mu$ as well as real $\mu$ using the thermodynamic geometry. This paper is organized as follows. In Sec. II, the RW periodicity and transition are briefly reviewed. In Sec. III, we show our formulation of the HRG model with EVE. In Sec. IV, we briefly review the thermodynamic geometry. In Sec. V, the numerical results of the scalar curvature and the phase structures are shown. In Sec. VI, we investigate the limiting temperature of our HRG model with EVE. Using the limiting temperature, an approximate location of the critical point and a simple empiric sufficient condition for quark deconfinement are obtained. The baryon radius dependence of the results is discussed in Sec. VII. Section VIII is devoted to the summary and discussions.
@@ II Roberge-Weiss periodicity and transition
${\cal L}(\theta_{\rm q})$ $=\bar{\psi}(\gamma_{\mu}D_{\mu}-m_{0})\psi-{1\over{4}}F_{\mu\nu}^{2}-i{\theta_{\rm q}\over{\beta}}\bar{\psi}\gamma_{4}\psi.$ (5)
Here $\psi$, $A_{\mu}$, $F_{\mu\nu}$, $D_{\mu}$, and $m_{0}$ are the quark field, the gluon field, the gluon field strength, the covariant derivative and the current quark mass matrix, respectively, and $\beta=1/T$. We use the Euclidean notation in Eqs. (3) $\sim$ (5).
@@
Next, we perform another transformation of the quark and gluon fields as follows:
@@
When $T<T_{\rm RW}$ where $T_{\rm RW}$ is the RW temperature, the net quark number density $n_{\rm Q}$ is a smooth function of $\theta_{\rm q}$. However, when $T>T_{\rm RW}$, it is discontinuous at $\theta_{\rm q}=(2k+1)\pi/3$ where $k$ is any integer, due to the degeneracy of the ground state. This discontinuity is called the RW transition. Similarly, the pressure $P$ of the system is a smooth function of $\theta_{\rm q}$ when $T<T_{\rm RW}$, but it has a cusp at $\theta_{\rm q}=(2k+1)\pi/3$ when $T>T_{\rm RW}$. The RW periodicity and the RW transition are confirmed by LQCD simulations de Forcrand and Philipsen (2002b); D’Elia and Lombardo (2003b) and $T_{\rm RW}$ is estimated as $195(\pm 1)\sim 208(\pm 5)$ MeV for the 2+1 flavor LQCD simulation Bonati et al. (2016); Cuteri et al. (2022); Bonati et al. (2019).
@@ III Hadron resonance gas model with excluded volume effects
where $P_{\rm M}$ is the pressure of mesons. For the mesons, we use an ideal Bose gas approximation. The other thermodynamic quantities are also calculated by using the thermodynamic relations. Our formalism of the HRG model with EVE is different from that of Refs. Taradiy et al. (2019); Savchuk et al. (2020); Vovchenko et al. (2017) . Comparison between two models is shown in Refs. Oshima et al. (2025) and our model is consistent with the model in Ref. Taradiy et al. (2019); Savchuk et al. (2020); Vovchenko et al. (2017) at least below $T_{\rm RWL}$.
@@
Hence, $n_{\rm B}$ is divergent at $\theta=(2n+1)\pi$ for any integer $n$ as is seen in Fig. 1. This singularity is nothing but the RW-like singularity which occurs at $T=T_{\rm RWL}=0.2103$ GeV. The value 0.2103 GeV is slightly smaller than the value 0.2108 GeV in Ref. Oshima et al. (2025), since we have updated the data of hadron resonances Navas and others (2024). On the other hand, as was already shown in Ref. Oshima et al. (2025), the use of the Boltzmann distribution approximation hardly changes the value of $T_{\rm RWL}$. If we use the nucleon gas model instead of the baryon resonances gas model, we obtain $T_{\rm RWL}=0.316$ GeV which is much higher than $T_{\rm RW}\sim 0.2$ GeV. The effects of the baryon resonances are very important for the realization $T_{\rm RWL}\sim T_{\rm RW}$.
@@
Comparing Fig. 1 (Fig. 4) with Fig. 2 (Fig. 3), we see that the saturation (linearization) of $n_{\mathrm{B}}$ ($P$) in the limit $\mu\to\pm\infty$ is related to the RW-like singularity at $\theta=(2n+1)\pi$. This indicates that the properties of the RW-like singularity in the imaginary chemical potential region contains significant information of the interaction among baryons in the large real chemical potential region.
For comparison, we show the $\theta_{\rm q}(=\theta/3)$-dependence of the number density and pressure in the PNJL model with imaginary chemical potential. The PNJL model Fukushima (2004); Ratti et al. (2006); Ghosh et al. (2006); Megias et al. (2006); Rößner et al. (2007); Kashiwa et al. (2008) is one of the most successful effective models of QCD and can also reproduce several important features of QCD in the region of the imaginary chemical potential when $T$ is not so small; for example, see Ref. Kashiwa (2019) as a review.
Fig. 5 shows the $\theta_{\rm q}$-dependence of the net baryon number density $n_{\rm Q}/3$ obtained by the PNJL model, when $T=T_{\rm RWL}=0.2103$ GeV. We see that the net baryon number density is discontinuous at $\theta=\pm\pi/3$. (Note that, in this calculation, $T=T_{\rm RWL}=0.2103$ GeV is larger than $T_{\rm RW}=0.201~{\rm GeV}$. Also note that the net baryon (quark) number density is pure imaginary in the region of the imaginary $\mu~(\mu_{q})$. )
@@
It seems that $T_{\rm RWL}$ is slightly higher than $T_{\rm RW}$. Hence, our scenario of the hadron-quark transition at $\theta=(2n+1)\pi$ is as follows. As $T$ approaches $T_{\rm RWL}$, the baryons become unstable and transformation from baryon matter to quark matter takes place when $T$ is equal to the temperature which is slightly lower than $T_{\rm RWL}$. This transition temperature is nothing but the Roberge-Weiss temperature $T_{\rm RW}$. However, the nature of the interaction which induces the instability is not clear yet. Hence, in this paper, we will investigate the nature of the interaction using the thermodynamic geometry.
@@ IV Thermodynamic geometry
The scalar curvature $R$ can be used to investigate the nature of the interaction. It is expected that the interaction in the system is repulsive (attractive), when $R>0$ ($R<0$). Using this criterion, it can be concluded that the ideal fermion gas has the repulsive nature, while the ideal boson gas exhibits an attractive one Janyszek and Mrugaa (1990). Hence, the interaction in baryon (meson) gas is expected to be repulsive (attractive).
@@
The thermodynamic geometry can be also used to determine the phase diagram. In $\mu$-$T$ plane, the curve on which $R=0$ may be the transition line of the phase transition. Castorina, Imbrosciano and Lanteri used the thermodynamic geometry to investigate the QCD phase diagram in $\mu$-$T$ plane Castorina et al. (2018). Since there is a sign problem in finite density LQCD, they used the power series expansion method to estimate $\phi$ and calculated the scalar curvature $R$. They used the $R=0$ criterion to determine the pseudo-critical temperature of the crossover transition. The obtained pseudo-critical temperature is consistent with that obtained by using the standard criterion in the LQCD. They also used the HRG model to estimate $\phi$ and calculated $R$. They found that the pseudo-critical temperature obtained in the HRG model is somewhat higher than that in LQCD. In Ref. Castorina et al. (2019), they performed the similar analyses but used also the HRG model with EVE. In this case, they found that the pseudo-critical temperature obtained by the HRG model with EVE is consistent with that in LQCD.
@@
The scalar curvature $R$ has the same dimension as the spatial volume. Hence, it is expected that, in proximity of a second order phase transition, $R$ is proportional to $\xi^{3}$ where $\xi$ is the correlation length and diverges when the second order phase transition takes place. Castorina, Lanteri and Mancani  Castorina et al. (2020a) studied the thermodynamic geometry in Nambu–Jona-Lasinio model. They found that the scalar curvature $R$ shows the divergent behavior at the critical temperature of the chiral phase transition when the current quark mass is zero. They also found that $R$ becomes large when the system approaches the critical point when the current quark mass is finite. Zhang, Wan and Ruggieri did the similar studies by using the Quark-meson model Zhang et al. (2020). Castorina, Lanteri and Ruggieri studied the fluctuations and the thermodynamic geometry of the chiral phase transition in the Quark-meson model Castorina et al. (2020b). Murgana, Greco, Ruggieri and Zappalà studied the thermodynamic geometry of the Quark-meson model using the functional renormalization group Murgana et al. (2024).
@@
At a glance, the two criteria, namely, the $R=0$ criterion and the $|R|=\infty$ criterion, contradict each other. However, in many cases, the divergence of $R$ is accompanied by the change of sign of $R$ Murgana et al. (2024). In particular, the sign change of $R$ takes place in the vicinity of the critical point where $R$ shows the divergent behavior Murgana et al. (2024). As will be seen later, there is no divergence of $R$ in the HRG model when the chemical potential $\mu$ is real. Hence, we mainly use the $R=0$ criterion to investigate the phase structure in the HRG model as was in the previous studies Castorina et al. (2018, 2019).